OpenStudy (anonymous):

Write the equation of the circle containing the points J(-3, -3), K(1, -3), and L(1, 1)

OpenStudy (badhi):

general form of the circle, $$x^2+y^2+cx+dy+e=0$$ since J,K,L go through the circle, $$(-3)^2+(-3)^2+c(-3)+d(-3)+e=0\\ 3c+3d-e=18\qquad(1)$$ $$(1)^2+(-3)^2+c(1)+d(-3)+e=0\\ 3d-c-e=10\qquad(2)$$ $$(1)^2+(1)^2+c(1)+d(1)+e=0\\ c+d+e=-2\qquad(3)$$ 3x(3)-(1) $$4e=-24\implies e=-6$$ (2)+(3) $$4d=8\implies d=2$$ $$c=2$$ equation of the circle- $$x^2+y^2+2x+2y=6$$

OpenStudy (anonymous):

thanks